# Go with Your Gut: Scaling Confidence for Autoregressive Image Generation

### v1

@@ 1 Introduction
Large language models (LLMs) (Brown et al., 2020; Vaswani et al., 2017; Radford et al., 2019) have demonstrated the capabilities of next-token prediction (NTP) paradigm. This success has renewed interest in applying autoregressive (AR) architectures beyond text, motivating recent visual generative models that represent images in discrete token spaces (Sun et al., 2024; Tian et al., 2024; Li et al., 2024) as shown in Figure 2 (b). Compared to diffusion models, which operate over continuous noise trajectories, token-based AR models promise a more unified modality interface.
As the field evolves, the parameters and training data of foundation models (Wang et al., 2024; Yang et al., 2025) have increasingly grown to levels that are inaccessible for most university researchers. In this context, many studies have started to investigate post-training methods. Inspired by recent advancements such as GRPO (Shao et al., 2024), a surge of reinforcement learning research has emerged in both language and visual generation domains (Jiang et al., 2025; Cui et al., 2025). Meanwhile, another research avenue focusing on test-time scaling (TTS) has emerged (Lightman et al., 2023; Muennighoff et al., 2025; Zuo et al., 2025), which aims to explore whether a slight increase in computational expense during inference can achieve performance on par with training-time methods, which typically incur much larger costs.
While test-time scaling has been extensively researched in language models, analogous progress for autoregressive visual generation remains sparse. Images differ from text in three practical ways that complicate direct transfer: (i) holism: dropping the last $20\%$ of a text sequence may still leave a syntactically valid answer, whereas truncating an image token stream yields an unusable artifact; (ii) objective ambiguity: many language scaling setups optimize toward a verifiable final answer (e.g., math reasoning), whereas image generation lacks a single ground-truth target; and (iii) early signal scarcity: partial image token decodes are visually unstable, making premature selection risky. Moreover, recent work TTS-VAR (Chen et al., 2025c) introduced TTS for the next-scale prediction (NSP) paradigm in visual autoregressive model (VAR) (Tian et al., 2024) by predicting images in a coarse-to-fine manner (Figure 2 (a)). This intermediate visibility enables reward models to score during scaling but comes with limitations that require predicting large residual token maps at each scale and frequent decoding makes the process inefficient and less suitable for the NTP paradigm.
@@ 2 Related Work
Autoregressive models have leveraged the scaling capabilities of language models (Yang et al., 2025; Brown et al., 2020; Radford et al., 2019) to generate images. These approaches employ discrete image tokenizers (Van Den Oord et al., 2017; Razavi et al., 2019) in conjunction with transformers, using a next-token prediction strategy. VQ-based methods (Lee et al., 2022; Razavi et al., 2019; Esser et al., 2021), e.g., VQ-VAE (Van Den Oord et al., 2017), convert image patches into index-based tokens, which are then predicted sequentially by a decoder-only transformer. However, these VQ-based AR methods are limited by the lack of scaled-up transformers and the inherent quantization error in VQ-VAE. This has prevented them from achieving performance on par with diffusion models. Recent advancements (Wu et al., 2025a; Yu et al., 2022; Team, 2024) have scaled up AR models for visual generation. Additionally, some variants have been proposed, such as the next-scale prediction paradigm of VAR (Tian et al., 2024; Han et al., 2025), which predicts from coarse to fine token maps, and the parallel token prediction of masked AR (MAR) (Li et al., 2024; Wu et al., 2025b; Fan et al., 2025). Despite these developments, the mainstream approach remains the NTP paradigm, particularly as the field moves towards unified models (Xie et al., 2025; Wang et al., 2024; Ge et al., 2024) that can jointly handle textual and visual tokens. This alignment with language modeling allows for more versatile and scalable architectures.
@@
Current LLMs have increasingly succeeded by allocating substantial reasoning at inference time, a paradigm known as test-time scaling (Snell et al., 2024; Welleck et al., 2024). This scaling can occur along two main axes: (1) Chain-of-Thought (CoT) (Wei et al., 2022) Depth: lengthening a single reasoning trajectory through more thinking steps, often relying on large-scale reinforcement learning with many samples (Yang et al., 2025; Jaech et al., 2024; Guo et al., 2025) or simpler post-training strategies (Ye et al., 2025; Muennighoff et al., 2025); (2) Parallel Generation: scaling by increasing the number of trajectories and aggregating them, as seen in works like Self-Consistency (Wang et al., 2023) and Best-of-N (Lightman et al., 2023). Recent efforts (Kang et al., 2025; Fu et al., 2025) have also integrated confidence estimation through token entropy into the test-time reasoning process, allowing the quality of individual traces to be assessed before aggregation with the rewards for majority voting (Wang et al., 2023). However, exploring TTS for AR image generation has been limited. This is due to the holistic nature of image generation, where overall coherence is paramount (see Figure 2 (c)), unlike reasoning tasks with well-defined ground truths. Additionally, the frequent early decoding required for images can be more computationally expensive than for language, suggesting that direct transfer of many LLM TTS techniques may not be suitable or optimal. To address this gap, we propose the first TTS strategy tailored for AR image generation. Notably, we pioneer the exploration of token entropy in image generation, enabling our method to leverage visual token confidence without the need for early decoding or additional rewards.
@@ 3 Preliminaries
Token entropy is a critical metric for evaluating the uncertainty associated with the predictions made by language models (Kang et al., 2025). It quantifies the amount of unpredictability in the model’s output distribution for a given token. The entropy $H$ at a specific position $i$ in the sequence can be defined as:
@@
Furthermore, token confidence can be derived from the predicted distribution (Fu et al., 2025). The confidence $C_{i}$ for a token at position $i$ is defined as:
@@ 4.1.2 Conditional Channel: Text Utilization Strength
$D_{t}=0.5+0.5\frac{K_{t}^{\mathrm{clip}}}{z_{\mathrm{max}}}.$ (10)
@@ 4.2.2 Guidance Scheduler
Table 1: Evaluation on GenEval (Ghosh et al., 2023) and TIIF-Bench (Wei et al., 2025) benchmarks. “Diff.+AR” refers to the unified architecture, and “MAR” indicates the masked AR architecture (Li et al., 2024). We bold the best results, and $\uparrow$” denotes that higher is better.
@@
  DALLE·3 (Betker et al., 2023) - Diff. - - - 0.67 78.40 68.45 62.69 72.94
Show-o (Xie et al., 2025) 1.3B Diff.+AR 0.80 0.31 0.50 0.68 71.30 59.89 68.66 59.24
LightGen (Wu et al., 2025b) 0.8B MAR 0.65 0.22 0.43 0.62 53.99 45.76 59.70 46.42
Infinity (Han et al., 2025) 2B VAR 0.85 0.49 0.57 0.73 71.63 57.81 61.19 59.66
Emu3 (Han et al., 2025) 8.5B AR 0.81 0.49 0.45 0.66 - - - -
Janus (Wu et al., 2025a) 1.5B AR 0.68 0.46 0.42 0.61 - - - -
AR-GRPO (Yuan et al., 2025) 0.8B AR 0.27 0.02 0.03 0.31 19.59 14.91 17.91 16.22
@@ 5.1 Experimental Settings
We apply ScalingAR to the advanced models: LlamaGen ($512\times 512$) (Sun et al., 2024) and AR-GRPO ($256\times 256$) (Yuan et al., 2025). Since no prior work has explored TTS for the NTP image generation, we focus our comparisons on the following conventional baselines: Importance Sampling (IS) (Owen & Zhou, 2000) and Best-of-N (BoN) (Lightman et al., 2023). We also provide results from Show-o (Xie et al., 2025), LightGen (Wu et al., 2025b), Infinity (Han et al., 2025), Emu3 (Wang et al., 2024), Janus (Wu et al., 2025a), and DALLE·3 (Betker et al., 2023) for reference.
@@
To evaluate the effectiveness of ScalingAR, we adopt GenEval (Ghosh et al., 2023) and TIIF-Bench (Wei et al., 2025) as primary benchmarks for both general and compositional text-to-image generation capabilities. These benchmarks offer a comprehensive evaluation of the model’s ability to produce high-quality and semantically consistent images from text prompts.
@@ 5.4 Scalability & Robustness Analysis
To answer RQ4, we compare ScalingAR with other TTS strategies (i.e., IS and BoN) in scaling width (i.e., sample number $N$) and depth (i.e., token length), as shown in Figure 5 (Right). To further assess the robustness of ScalingAR, we adopt the idea of “impossible prompting” (Bai et al., 2025b; Chen et al., 2025a) (e.g., “A young boy … using chopsticks as a writing instrument, … in a photo-realistic…”) to evaluate its performance even when none of the candidates are ideal, with the results presented in Figure 6. Our observations are summarized as follows: Obs.ScalingAR unlocks scalable generalization across both width and depth. As shown in Figure 5 (Right), ScalingAR consistently outperforms IS and BoN across varying sample numbers and token lengths. This suggests that our scaling strategy enables performance to scale up effectively as scaling width and depth increase, making it a reliable solution for diverse autoregressive tasks. Obs.ScalingAR empowers robust generation beyond standard scenarios. Figure 6 (Left) demonstrates that under impossible prompts for unrealistic scenarios, ScalingAR exhibits clear robustness advantages over baselines. Figure 6 (Right) further confirms that our method achieves more effective scaling when generating under challenging conditions, highlighting its adaptability and reliability in adverse scenarios.
@@ A.1 More Details of Experimental Settings
To evaluate the robustness of ScalingAR, we further employ prompts from IPV-Txt from Impossible Videos [ICML’25] (Bai et al., 2025b). Specifically, since IPV-Txt was originally designed to assess the capabilities of video understanding (Wang et al., 2025; Huang et al., 2025; Bai et al., 2025a) and video generation models (Chen et al., 2025b; Wan et al., 2025; Shao et al., 2025), we filtered prompts suitable for image generation from IPV-Txt, then employed Impossible Prompt Following (IPF) as the evaluation metric, which measures the alignment between generated images and the semantic intent of impossible prompts. Following Bai et al. (2025b), we employed GPT-4o to perform binary judgments on each image based on prompt adherence. For qualitative results in Figure 6 (Right):
@@ Appendix D The Use of LLMs
Figure 9: More results demonstrations of ScalingAR on AR-GRPO (Yuan et al., 2025).
Figure 10: More results demonstrations of ScalingAR on LlamaGen (Sun et al., 2024).

### v2

@@ 1 Introduction
Large language models (LLMs) (brown2020language; vaswani2017attention; radford2019language) have demonstrated the capabilities of next-token prediction (NTP) paradigm. This success has renewed interest in applying autoregressive (AR) architectures beyond text, motivating recent visual generative models that represent images in discrete token spaces (sun2024autoregressive; tian2024visual; li2024autoregressive) as shown in Figure 2 (b). Compared to diffusion models, which operate over continuous noise trajectories, token-based AR models promise a more unified modality interface.
As the field evolves, the parameters and training data of foundation models (wang2024emu3; qwen3) have increasingly grown to levels that are inaccessible for most university researchers. In this context, many studies have started to investigate post-training methods. Inspired by recent advancements such as GRPO (shao2024deepseekmath), a surge of reinforcement learning research has emerged in both language and visual generation domains (jiang2025t2i; cui2025entropy). Meanwhile, another research avenue focusing on test-time scaling (TTS) has emerged (lightman2023let; muennighoff2025s1; zuo2025ttrl), which aims to explore whether a slight increase in computational expense during inference can achieve performance on par with training-time methods, which typically incur much larger costs.
While test-time scaling has been extensively researched in language models, analogous progress for autoregressive visual generation remains sparse. Images differ from text in three practical ways that complicate direct transfer: (i) holism: dropping the last $20\%$ of a text sequence may still leave a syntactically valid answer, whereas truncating an image token stream yields an unusable artifact; (ii) objective ambiguity: many language scaling setups optimize toward a verifiable final answer (e.g., math reasoning), whereas image generation lacks a single ground-truth target; and (iii) early signal scarcity: partial image token decodes are visually unstable, making premature selection risky. Moreover, recent work TTS-VAR (chen2025tts) introduced TTS for the next-scale prediction (NSP) paradigm in visual autoregressive model (VAR) (tian2024visual) by predicting images in a coarse-to-fine manner (Figure 2 (a)). This intermediate visibility enables reward models to score during scaling but comes with limitations that require predicting large residual token maps at each scale and frequent decoding makes the process inefficient and less suitable for the NTP paradigm.
@@ 2 Related Work
Autoregressive models have leveraged the scaling capabilities of language models (qwen3; brown2020language; radford2019language) to generate images. These approaches employ discrete image tokenizers (van2017neural; razavi2019generating) in conjunction with transformers, using a next-token prediction strategy. VQ-based methods (lee2022autoregressive; razavi2019generating; esser2021taming), e.g., VQ-VAE (van2017neural), convert image patches into index-based tokens, which are then predicted sequentially by a decoder-only transformer. However, these VQ-based AR methods are limited by the lack of scaled-up transformers and the inherent quantization error in VQ-VAE. This has prevented them from achieving performance on par with diffusion models. Recent advancements (wu2025janus; yu2022scaling; team2024chameleon) have scaled up AR models for visual generation. Additionally, some variants have been proposed, such as the next-scale prediction paradigm of VAR (tian2024visual; han2025infinity), which predicts from coarse to fine token maps, and the parallel token prediction of masked AR (MAR) (li2024autoregressive; wu2025lightgen; fan2025fluid). Despite these developments, the mainstream approach remains the NTP paradigm, particularly as the field moves towards unified models (xie2025showo; wang2024emu3; ge2024seed) that can jointly handle textual and visual tokens. This alignment with language modeling allows for more versatile and scalable architectures.
@@
Current LLMs have increasingly succeeded by allocating substantial reasoning at inference time, a paradigm known as test-time scaling (snell2024scaling; welleck2024from). This scaling can occur along two main axes: (1) Chain-of-Thought (CoT) (wei2022chain) Depth: lengthening a single reasoning trajectory through more thinking steps, often relying on large-scale reinforcement learning with many samples (qwen3; jaech2024openai; guo2025deepseek) or simpler post-training strategies (ye2025limo; muennighoff2025s1); (2) Parallel Generation: scaling by increasing the number of trajectories and aggregating them, as seen in works like Self-Consistency (wang2023selfconsistency) and Best-of-N (lightman2023let). Recent advances (kang2025scalable; fu2025deep) have also incorporated token entropy for confidence estimation, improving the quality of reasoning traces. However, exploring TTS for AR image generation has been limited. This is due to the holistic nature of image generation, where overall coherence is paramount (see Figure 2 (c)), unlike reasoning tasks with well-defined ground truths. Recent work guo2025can applies Best-of-N to AR image generation by introducing external reward models but without new TTS strategy design itself. Moreover, similar to TTS-VAR (chen2025tts), it relies on frequent early decoding and rewards. To this end, we propose the first TTS strategy tailored for AR image generation. Notably, we pioneer the exploration of token entropy in image generation, enabling our method to leverage visual token confidence without the need for early decoding or additional rewards.
@@ 3 Preliminaries
Token entropy is a critical metric for evaluating the uncertainty associated with the predictions made by language models (kang2025scalable). It quantifies the amount of unpredictability in the model’s output distribution for a given token. The entropy $H$ at a specific position $i$ in the sequence can be defined as:
@@
Furthermore, token confidence can be derived from the predicted distribution (fu2025deep). The confidence $C_{i}$ for a token at position $i$ is defined as:
@@ 4.1.2 Conditional Channel: Text Utilization Strength
$\widehat{D}_{t}=0.5+0.5\frac{K_{t}^{\mathrm{clip}}}{z_{\mathrm{max}}}.$ (10)
@@ 4.2.2 Guidance Scheduler
Table 1: Evaluation on GenEval (ghosh2023geneval) and TIIF-Bench (wei2025tiif) benchmarks. “Diff.+AR” refers to the unified architecture, and “MAR” indicates the masked AR architecture (li2024autoregressive). We bold the best results, and $\uparrow$” denotes that higher is better.
@@
  DALLE·3 (betker2023improving) - Diff. - - - 0.67 78.40 68.45 62.69 72.94
Show-o (xie2025showo) 1.3B Diff.+AR 0.80 0.31 0.50 0.68 71.30 59.89 68.66 59.24
LightGen (wu2025lightgen) 0.8B MAR 0.65 0.22 0.43 0.62 53.99 45.76 59.70 46.42
Infinity (han2025infinity) 2B VAR 0.85 0.49 0.57 0.73 71.63 57.81 61.19 59.66
Emu3 (han2025infinity) 8.5B AR 0.81 0.49 0.45 0.66 - - - -
Janus (wu2025janus) 1.5B AR 0.68 0.46 0.42 0.61 - - - -
AR-GRPO (yuan2025argrpo) 0.8B AR 0.27 0.02 0.03 0.31 19.59 14.91 17.91 16.22
@@ 5.1 Experimental Settings
We apply ScalingAR to the advanced models: LlamaGen ($512\times 512$) (sun2024autoregressive) and AR-GRPO ($256\times 256$) (yuan2025argrpo). Since no prior work has explored TTS for the NTP image generation, we focus our comparisons on the following conventional baselines: Importance Sampling (IS) (owen2000safe) and Best-of-N (BoN) (lightman2023let). We also provide results from Show-o (xie2025showo), LightGen (wu2025lightgen), Infinity (han2025infinity), Emu3 (wang2024emu3), Janus (wu2025janus), and DALLE·3 (betker2023improving) for reference.
@@
To evaluate the effectiveness of ScalingAR, we adopt GenEval (ghosh2023geneval) and TIIF-Bench (wei2025tiif) as primary benchmarks for both general and compositional text-to-image generation capabilities. These benchmarks offer a comprehensive evaluation of the model’s ability to produce high-quality and semantically consistent images from text prompts.
@@ 5.4 Scalability & Robustness Analysis
To answer RQ4, we compare ScalingAR with other TTS strategies (i.e., IS and BoN) in scaling width (i.e., sample number $N$) and depth (i.e., token length), as shown in Figure 5 (Right). To further assess the robustness of ScalingAR, we adopt the idea of “impossible prompting” (bai2025impossible; chen2025hierarchical) (e.g., “A young boy … using chopsticks as a writing instrument, … in a photo-realistic…”) to evaluate its performance even when none of the candidates are ideal, with the results presented in Figure 6. Our observations are summarized as follows: Obs.ScalingAR unlocks scalable generalization across both width and depth. As shown in Figure 5 (Right), ScalingAR consistently outperforms IS and BoN across varying sample numbers and token lengths. This suggests that our scaling strategy enables performance to scale up effectively as scaling width and depth increase, making it a reliable solution for diverse autoregressive tasks. Obs.ScalingAR empowers robust generation beyond standard scenarios. Figure 6 (Left) demonstrates that under impossible prompts for unrealistic scenarios, ScalingAR exhibits clear robustness advantages over baselines. Figure 6 (Right) further confirms that our method achieves more effective scaling when generating under challenging conditions, highlighting its adaptability and reliability in adverse scenarios.
@@ A.1 More Details of Experimental Settings
To evaluate the robustness of ScalingAR, we further employ prompts from IPV-Txt from Impossible Videos [ICML’25] (bai2025impossible). Specifically, since IPV-Txt was originally designed to assess the capabilities of video understanding (wang2025internvl3_5; huang2025mathcalvistamathcaldpo; Qwen2.5-VL) and video generation models (chen2025temporal; wan2025; shao2025finephys), we filtered prompts suitable for image generation from IPV-Txt, then employed Impossible Prompt Following (IPF) as the evaluation metric, which measures the alignment between generated images and the semantic intent of impossible prompts. Following bai2025impossible, we employed GPT-4o to perform binary judgments on each image based on prompt adherence. For qualitative results in Figure 6 (Right):
@@ Appendix D The Use of LLMs
Figure 9: More results demonstrations of ScalingAR on AR-GRPO (yuan2025argrpo).
Figure 10: More results demonstrations of ScalingAR on LlamaGen (sun2024autoregressive).